# Computational Modeling for Personalized Transcranial Electrical Stimulation: Theory, Tools, and Applications

### v1

@@ 4.1 Refining Forward Models through Segmentation, Conductivity, and Anisotropy Updates
A substantial discrepancy often exists between the electric fields predicted by computational models and those empirically measured (Fig. 3). This mismatch is largely attributable to individual variability in tissue properties and limitations in anatomical segmentation. Improving segmentation fidelity, both in terms of accuracy and anatomical detail, is crucial for enhancing forward modeling precision [79, 143] (Fig. 3 B). For instance, Weise et al. [144] demonstrated that the inclusion of the meninges, often omitted due to computational cost, can significantly alter the peak electric field intensity. Similarly, Wang et al. [145] reported that layered versus single-layer individualized skull models yield markedly different field distributions. These findings underscore the importance of incorporating detailed and individualized tissue definitions into the model. This challenge is further compounded in pediatric or neonatal populations, where immature tissue development and limitations in imaging resolution complicate segmentation and conductivity estimation [146].
@@
Tissue conductivity is another key issue in modeling electric field distribution [2]. Conductivity values can vary significantly across individuals, tissue temperature, and even stimulation frequency [147]. To accurately set the tissue conductivity values, head model parameters can be calibrated using intracranial recordings, such as intracranial EEG (iEEG), by aligning predicted and measured values in localized brain regions (Fig.3 C). Several techniques have been proposed for estimating and optimizing conductivity, as well as for quantifying the reliability of model predictions[2, 148, 149]. For example, Huang et al. [2] developed an optimization algorithm that adjusts conductivity values to minimize the discrepancy between modeled and observed iEEG potentials. This calibration can enhance field estimation accuracy across the brain.
@@
Most conventional head models simplify conductivity as a scalar value, without considering the anisotropy of conductivity. However, this assumption overlooks the directional complexity of neural tissue. The presence of aligned white matter tracts gives rise to anisotropic conductivity, which significantly influences the spatial distribution of the electric field [150, 151] (Fig.3 D). Suh et al. [152] demonstrated that incorporating anisotropic skull conductivity can reduce off-target electric fields by 12-14%, thus enhancing spatial specificity. A commonly used approach for modeling anisotropy involves constructing conductivity tensors based on diffusion MRI-derived diffusion tensors and fractional anisotropy metrics. These tensors are rescaled according to known tissue conductivity values and integrated into FEM frameworks. This enables more biologically realistic simulations that account for directional dependencies in current propagation.
@@
Figure 3: The influence of model updates on the norm of the electrical field. (A) The Standard Model: The upper panel displays a segmentation with five tissue types, each assigned conductivities based on standard values from prior research (WM: 0.126; GM: 0.275; CSF: 1.654; Bone: 0.01; Scalp: 0.465). The lower panel illustrates the distribution of the electric field’s magnitude under a 1 mA tDCS from F3 to F4 in the EEG10-10 system. Following visualizations follow the same settings. (B) Bone subdivisions. The bone is segmented into two distinct categories: compact bone with a conductivity of 0.008 and spongy bone with a conductivity of 0.025. (C) Conductivity optimization. The conductivity of Gray Matter is adjusted to its highest reported value in previous research, being 0.6. In practice, conductivity optimization can be combined with sEEG to fine-tune the model’s parameters. (D) Anisotropy model. The upper panel depicts the conductivity of White Matter. The conductivity of tissues is configured as anisotropic by direct mapping based on linear rescaling of tensors from diffusion MRI. The upper threshold for conductivity is established at 2 S/m, with a maximum ratio of 10 between the largest and smallest conductivity eigenvalues. The unit of conductivity: S/m. The unit of electrical field: V/m. Abbr: WM, White Matter; GM, Gray Matter; CSF, Cerebrospinal Fluid; EF, electrical field; tDCS, transcranial Direct Current Stimulation.
@@ 4.2 Incorporating Brain-State Dynamics into Inverse Optimization
Conventional inverse optimization approaches in tES often treat the brain as a passive electrical medium, neglecting its intrinsic nonlinear and state-dependent behavior. However, accumulating evidence suggests that the efficacy of neuromodulation is highly dependent on the brain’s dynamic state at the time of stimulation [1, 153]. These brain states—reflected in EEG rhythms, BOLD fluctuations, or patterns of functional connectivity—can critically influence the brain’s responsiveness to external stimulation. Accordingly, integrating dynamic brain-state information into the optimization framework represents a key step toward adaptive and truly personalized tES.
@@
The first step in brain-state-informed optimization is the acquisition of relevant neural signals during tES. EEG is the most widely used modality to assess tES-induced neural changes before and after stimulation [154, 155, 156]. However, accurately capturing brain activity during stimulation remains challenging due to strong electrical artifacts. To address this, alternative modalities such as fMRI and sEEG have been employed for simultaneous recording with tES [156, 157, 158, 159, 160]. These studies have not only demonstrated the efficacy of tES but also revealed significant discrepancies between simulated and empirically observed neural responses [91, 159], highlighting the influence of individual variability and dynamic brain states. One explanation for these discrepancies is the oversimplification of the brain as a linear conductor in computational models. In reality, the brain is a complex, nonlinear, and adaptive system. Recent studies have shown that tES can modulate brain networks rather than isolated regions [71, 161, 162]. For example, Rostami et al. demonstrated that tACS modulates cortical connectivity within the dorsal attention network, as evidenced by changes in phase-locking values [163]. Mulyana et al. further advanced this direction by proposing a closed-loop tES-fMRI framework to optimize individualized stimulation of frontoparietal networks, resulting in enhanced functional connectivity and improved working memory [156]. Krause et al. suggested that individual variability in neuronal phase preference contributes to heterogeneous tES outcomes, which could be better explained using oscillator-based models [1, 164].
To better capture the temporal evolution of brain activity, hidden Markov models (HMMs) have been widely employed. For example, Kasten et al. used HMMs to analyze MEG data collected before and after tACS, revealing that spontaneous brain states and their underlying functional networks vary in susceptibility to stimulation [165]. Extending this approach, Brown et al. applied high-definition alpha-frequency tACS during a continuous performance task while recording fMRI data. Their results demonstrated that stimulation modulated the temporal dynamics of brain states, as identified by HMMs [166]. We can design objective functions based on brain states and obtain corresponding stimulation strategies. Combining the method with control theory, the brain can be modeled as a dynamic system influenced by external input. A simplified representation is given by a discrete-time linear time-invariant system:
@@
Control strategies can be broadly categorized as either model-driven or data-driven. In model-driven control, matrix $A$ is usually derived from prior knowledge, such as structural (white matter tracts), functional (statistical dependencies), or effective (causal interactions) connectivity [167, 168, 169]. The leadfield matrix from electrical modeling can serve as matrix $B$, quantifying how each electrode influences brain regions. In contrast, data-driven approaches aim to estimate both $A$ and $B$ directly from empirical neural recordings. Given time-series data that reflect evolving brain states $x(t)$, the goal is to design optimal control inputs $u(t)$ that drive the system toward desirable cognitive or clinical outcomes.
@@ 5 Limitations and future directions
To fully harness the advantages of computational models, their integration with real-world stimulation hardware is essential. However, this integration is complicated by the diversity of stimulation devices and experimental environments. Most models are built on idealized assumptions, such as standardized equipment configurations and accurate parameter settings, which may not reflect real-world conditions. In practice, factors such as head curvature and subject motion during experiments can result in uneven contact pressure between electrodes and the scalp. This may displace electrodes or alter contact quality, ultimately distorting the electric field distribution and reducing stimulation efficacy [170]. Although methods have been developed for real-time electrode localization [171, 172], current tES frameworks generally lack the capability to dynamically account for such variabilities and adapt the model accordingly. Recent studies have begun to explore more adaptive and online modeling strategies to address this limitation [173, 174]. Additionally, variability in electrode properties—such as size, material, and surface texture—can significantly influence stimulation outcomes [175, 176]. Despite its importance, incorporating these properties into optimization algorithms remains challenging, largely due to the absence of well-defined models to quantify their effects. While electrode size has been shown to play a critical role, no universally accepted equations currently exist to capture the influence of other electrode characteristics, such as material composition and geometry.
@@
The analysis of brain signals during tES presents a promising direction, particularly with the emergence of tES devices that support simultaneous neural signal acquisition, thereby enabling the development of closed-loop systems. Among these, EEG is frequently employed alongside tES due to the compatibility of their underlying principles. However, recording EEG during ongoing tES poses significant technical challenges. The stimulation introduces substantial artifacts, often several orders of magnitude larger than the neural signals of interest, and may even saturate EEG amplifiers [3, 177]. In addition to electrical artifacts, physiological noise from sources such as cardiac and respiratory activity further contaminates the EEG recordings [178, 177]. As a result, the signal-to-noise ratio (SNR) is drastically reduced, making direct interpretation of EEG during tES infeasible. To mitigate this issue, denoising techniques have been proposed. Several offline denoising algorithms have been developed for tES-EEG data [179, 180, 181], and more recently, real-time approaches have emerged to facilitate online monitoring of brain activity. For example, Guarnieri et al. introduced an adaptive spatial filtering method, known as alternating current regression (AC-REG), specifically designed to attenuate tACS-induced artifacts in real time [182]. Despite these advances, major challenges remain. The most important of these is the lack of ground truth data, which makes it difficult to verify whether the denoised EEG accurately reflects true neural activity or still contains residual artifacts. Moreover, since tES itself modulates brain activity, the denoising process risks removing not only artifacts but also genuine stimulation-induced neural changes. As a result, validating the effectiveness and fidelity of denoised signals remains an open and actively investigated issue [183].
@@
The frequency of the injected current is another critical factor that warrants careful consideration in computational modeling. Contemporary models predominantly emphasize the amplitude of stimulation and the resulting electric field distribution, often overlooking the significant influence of current frequency. In the case of tACS, which typically operates at low frequencies, the quasi-static approximation is commonly employed. This approximation models the electric field at peak current amplitudes, under the assumption that temporal variations in current intensity only scale the magnitude of the electric field without affecting its spatial distribution. Nevertheless, an increasing number of studies have reported that frequency and phase can substantially affect both the distribution and efficacy of stimulation [67, 145]. These effects become especially pronounced at the higher carrier frequencies used in tTIS. For instance, Liu et al. observed a slight but measurable decrease in the field magnitude at temporal interference frequencies exceeding 2 kHz, which may be attributable to frequency-dependent changes in tissue conductivity [184, 185]. Beyond physical field properties, frequency also plays a critical role in shaping neural responses. Variations in stimulation frequency can significantly modulate neuronal excitability and synaptic dynamics, influencing the effectiveness of neuromodulation [186]. Therefore, it is imperative to expand computational models to incorporate frequency-dependent physiological responses. This can be achieved by integrating large-scale concurrent neural recordings and leveraging advanced signal analysis methods. Such data-driven approaches are essential for developing predictive models that elucidate the complex relationship between stimulation frequency and brain function.
@@
Additionally, the orientation of the electrical field is pivotal, as it can dictate the excitatory or inhibitory responses of neurons. Especially for tTIS, which regulated brain by two pairs of electrodes, a given field orientation can significantly enhance the focality of the envelope field. While optimization algorithms have been developed to refine this aspect, there remains a gap in our understanding of how to precisely configure the field’s direction for specific targets. In the case of the cerebral cortex, a vertical orientation is commonly employed [60]. However, for other tissues, defining a more precise target necessitates a deeper integration of neuroscience knowledge. Designing the objective function based on advanced neuroimaging could be a viable approach. For instance, fiber orientation obtained through diffusion MRI could be incorporated into the objective function.

### v2

@@ 4.1 Refining Forward Models through Segmentation, Conductivity, and Anisotropy Updates
A substantial discrepancy often exists between the electric fields predicted by computational models and those empirically measured (Fig. 3). This mismatch is largely attributable to individual variability in tissue properties and limitations in anatomical segmentation. Improving segmentation fidelity, both in terms of accuracy and anatomical detail, is crucial for enhancing forward modeling precision [puonti2020accurate, jalal2021robust] (Fig. 3 B). For instance, Weise et al. [weise2022effect] demonstrated that the inclusion of the meninges, often omitted due to computational cost, can significantly alter the peak electric field intensity. Using a boundary element-fast multipole method to incorporate the meninges into electric field simulations, they found that the cortical electric field strength increased by approximately 30% during tES when the meninges were included. Similarly, Wang et al. [wang2021influence] reported that in tTIS, the Pearson correlation coefficient between electric field distributions derived from layered-skull and single-layer skull models was only 0.746, which highlights substantial discrepancies in predicted electric field distributions across forward models employing different skull representation strategies. These findings underscore the importance of incorporating detailed and individualized tissue definitions into the model. This challenge is further compounded in pediatric or neonatal populations, where immature tissue development and limitations in imaging resolution complicate segmentation and conductivity estimation [despotovic2013realistic].
Additionally, a recent trend in development is the integration of deep learning into head model segmentation. It has demonstrated outstanding performance in image segmentation and has been increasingly applied to the delineation of anatomical brain structures. Rashed et al. [rashed2019development] introduced ForkNet, which leverages convolutional neural networks (CNNs) for automatic tissue segmentation from MRI data, thereby enhancing both the accuracy and efficiency of head tissue delineation. Hirsch et al. [hirsch2021segmentation] developed MultiPrior, a model that integrates deep volumetric networks with multiple spatial priors to improve segmentation accuracy and robustness. Stolte et al. [stolte2024precise] introduced the GRACE model, which effectively segmented personalized head models of elderly individuals, facilitating high-fidelity modeling of age-related neurological conditions.
@@
Tissue conductivity is another key issue in modeling electric field distribution [huang2017measurements]. Conductivity values can vary significantly across individuals, tissue temperature, and even stimulation frequency [geddes1967specific]. To accurately set the tissue conductivity values, head model parameters can be calibrated using intracranial recordings, such as intracranial EEG (iEEG), by aligning predicted and measured values in localized brain regions (Fig.3 C). Several techniques have been proposed for estimating and optimizing conductivity, as well as for quantifying the reliability of model predictions. For example, Huang et al. [huang2017measurements] developed an optimization algorithm that adjusts conductivity values to minimize the discrepancy between modeled and observed iEEG potentials. Carvallo et al. [carvallo2018biophysical] introduced a biophysical modeling framework that utilizes stereoelectroencephalography (sEEG) data to estimate tissue-specific conductivity parameters by comparing simulated and recorded potential distributions, thereby significantly improving the accuracy of individualized conductivity estimation across brain tissues. Saturnino et al. [saturnino2019principled] developed a systematic uncertainty analysis framework to assess how variability in tissue conductance affects electric field computations, offering a theoretical foundation for optimizing conductance parameters and calibrating individualized head models. This calibration can enhance field estimation accuracy across the brain.
@@
Most conventional head models simplify conductivity as a scalar value, without considering the anisotropy of conductivity. However, this assumption overlooks the directional complexity of neural tissue. The presence of aligned white matter tracts gives rise to anisotropic conductivity, which significantly influences the spatial distribution of the electric field [Tuch2001ConductivityTM, Opitz2011HowTB] (Fig.3 D). Suh et al. [Suh2012InfluenceOA] demonstrated that incorporating anisotropic skull conductivity can reduce off-target electric fields by 12-14%, thus enhancing spatial specificity. A commonly used approach for modeling anisotropy involves constructing conductivity tensors based on diffusion MRI-derived diffusion tensors and fractional anisotropy metrics. These tensors are rescaled according to known tissue conductivity values and integrated into FEM frameworks. This enables more biologically realistic simulations that account for directional dependencies in current propagation.
@@
Figure 3: The influence of model updates on the norm of the electric field. (A) The Standard Model: The upper panel displays a segmentation with five tissue types, each assigned conductivities based on standard values from prior research (WM: 0.126; GM: 0.275; CSF: 1.654; Bone: 0.01; Scalp: 0.465). The lower panel illustrates the distribution of the electric field’s magnitude under a 1 mA tDCS from F3 to F4 in the EEG10-10 system. Following visualizations follow the same settings. (B) Bone subdivisions. The bone is segmented into two distinct categories: compact bone with a conductivity of 0.008 and spongy bone with a conductivity of 0.025. (C) Conductivity optimization. The conductivity of Gray Matter is adjusted to its highest reported value in previous research, being 0.6. In practice, conductivity optimization can be combined with sEEG to fine-tune the model’s parameters. (D) Anisotropy model. The upper panel depicts the conductivity of White Matter. The conductivity of tissues is configured as anisotropic by direct mapping based on linear rescaling of tensors from diffusion MRI. The upper threshold for conductivity is established at 2 S/m, with a maximum ratio of 10 between the largest and smallest conductivity eigenvalues. The unit of conductivity: S/m. The unit of electric field: V/m. Abbr: WM, White Matter; GM, Gray Matter; CSF, Cerebrospinal Fluid; EF, electric field; tDCS, transcranial Direct Current Stimulation.
4.2 Coupling from Electric Field Distribution to Neural Activity
Beyond improving physical-level accuracy, another promising direction is to incorporate individualized neural dynamics into the model [deco2011emerging, breakspear2017dynamic]. Neural dynamics shape how external stimulation modulates the brain’s functional responses. Therefore, incorporating neural population dynamics into individualized modeling is essential for enhancing the physiological realism of computational models [cottone2018new]. A widely used framework is the Neural Mass Model (NMM) [jansen1995electroencephalogram, david2003neural]. NMM describes the average activity of excitatory and inhibitory neuronal populations within cortical columns by employing a nonlinear mapping between the mean membrane potential and firing rate of neuron groups. Compared with single-cell models, the NMM offers a balance between biophysical interpretability and computational efficiency, capturing macroscale oscillations, rhythmic synchronization, and dynamic responses to external stimulation. Consequently, it is widely employed for generating and interpreting EEG, MEG, and fMRI signals.
Neural mass models represent neural populations as fundamental units, forming local dynamical modules based on input–output relationships between excitatory and inhibitory subpopulations. In personalized modeling, researchers commonly use structural connectivity derived from imaging data (e.g., DTI) to determine the coupling matrix between nodes in the NMM. Model parameters are subsequently optimized against individual EEG or fMRI data to reproduce subject-specific neural dynamics [babajani2006integrated, sanz2015mathematical, cakan2023neurolib]. This structure-function integrated modeling framework enables the NMM to serve as the dynamic core of an individualized head model, allowing for predictions of neural responses and state transitions under various stimulation patterns.
In tES research, integrating NMM with electric field modeling offers a novel framework to understand how electrical stimulation regulates neural dynamics. Traditional volumetric conductor models treat the brain as a linear conductor, focusing on the distribution of the electric field. However, these models overlook the nonlinear and state-dependent characteristics of neural responses under external electric fields. In contrast, NMMs capture the mean activity of excitatory and inhibitory populations within cortical regions by deriving population-level equations from spiking neuron networks. This framework enables biologically realistic simulations of how external stimulation modulates local oscillatory rhythms, interregional synchrony, and large-scale network coupling [ali2013transcranial, neuling2012good]. Kunze et al. [kunze2016transcranial] integrated FEM with NMM to simulate large-scale networks, demonstrating that tDCS modulates resting-state connectivity and network dynamics.
Embedding NMM within stimulated regions of personalized head models allows simulation of functional-level neurodynamic responses. Karimi et al. [karimi2025precision] developed an NMM-based framework that combines MRI/DWI-derived connectivity with electric field simulations, enabling the optimization of individualized stimulation parameters. This approach enables the optimization of stimulation parameters to target specific regions while precisely regulating the whole-brain dynamics. Future advancements lie in coupling NMM with multi-scale neuronal models. For instance, Chung et al. [chung2022key] integrated FEM with single-neuron models to uncover how tES effects depend on neuron type, field orientation, and connectivity. Integrating neural dynamics into personalized models can capture individual variability in cortical excitability, oscillatory states, and connectivity. This approach facilitates the accurate prediction of stimulation effects and the optimization of parameters tailored to individual neural states.
@@ 4.3 Incorporating Brain-State Dynamics into Inverse Optimization
Conventional inverse optimization approaches in tES often treat the brain as a passive electrical medium, neglecting its intrinsic nonlinear and state-dependent behavior. However, accumulating evidence suggests that the efficacy of neuromodulation is highly dependent on the brain’s dynamic state at the time of stimulation [Krause2023TranscranialES, Frhlich2010EndogenousEF]. These brain states—reflected in EEG rhythms, BOLD fluctuations, or patterns of functional connectivity—can critically influence the brain’s responsiveness to external stimulation. Accordingly, integrating dynamic brain-state information into the optimization framework represents a key step toward adaptive and truly personalized tES.
@@
The first step in brain-state-informed optimization is the acquisition of relevant neural signals during tES. EEG is the most widely used modality to assess tES-induced neural changes before and after stimulation [Wang2023TheEO, Tashiro2020ProbingEA, Mulyana2022OnlineCR]. However, accurately capturing brain activity during stimulation remains challenging due to strong electrical artifacts. To address this, alternative modalities such as fMRI and sEEG have been employed for simultaneous recording with tES [louviot2022transcranial, Mulyana2022OnlineCR, GhobadiAzbari2020fMRIAT, Mulyana2021OnlineCR, Lou2024ADF]. These studies have not only demonstrated the efficacy of tES but also revealed significant discrepancies between simulated and empirically observed neural responses [louviot2022transcranial, luo2025frequency], highlighting the influence of individual variability and dynamic brain states. One explanation for these discrepancies is the oversimplification of the brain as a linear conductor in computational models. In reality, the brain is a complex, nonlinear, and adaptive system. Recent studies have shown that tES can modulate brain networks rather than isolated regions [saturnino2021optimizing, Ruffini2014OptimizationOM, Ali2013TranscranialAC]. For example, Rostami et al. demonstrated that tACS modulates cortical connectivity within the dorsal attention network, as evidenced by changes in phase-locking values [Rostami20206HT]. Mulyana et al. further advanced this direction by proposing a closed-loop tES-fMRI framework to optimize individualized stimulation of frontoparietal networks, resulting in enhanced functional connectivity and improved working memory [Mulyana2022OnlineCR]. Krause et al. suggested that individual variability in neuronal phase preference contributes to heterogeneous tES outcomes, which could be better explained using oscillator-based models [Krause2023TranscranialES, Krause2019TranscranialAC].
To better capture the temporal evolution of brain activity, hidden Markov models (HMMs) have been widely employed. For example, Kasten et al. used HMMs to analyze MEG data collected before and after tACS, revealing that spontaneous brain states and their underlying functional networks vary in susceptibility to stimulation [Kasten2022TheHB]. Extending this approach, Brown et al. applied high-definition alpha-frequency tACS during a continuous performance task while recording fMRI data. Their results demonstrated that stimulation modulated the temporal dynamics of brain states, as identified by HMMs [Brown2023TranscranialSO]. We can design objective functions based on brain states and obtain corresponding stimulation strategies. Combining the method with control theory, the brain can be modeled as a dynamic system influenced by external input. A simplified representation is given by a discrete-time linear time-invariant system:
@@
Control strategies can be broadly categorized as either model-driven or data-driven. In model-driven control, matrix $A$ is usually derived from prior knowledge, such as structural (white matter tracts), functional (statistical dependencies), or effective (causal interactions) connectivity [gu2015controllability, tanner2024multi, luo2025mapping]. The leadfield matrix from electrical modeling can serve as matrix $B$, quantifying how each electrode influences brain regions. In contrast, data-driven approaches aim to estimate both $A$ and $B$ directly from empirical neural recordings. Given time-series data that reflect evolving brain states $x(t)$, the goal is to design optimal control inputs $u(t)$ that drive the system toward desirable cognitive or clinical outcomes.
@@ 5 Limitations and Future Directions
To fully harness the advantages of computational models, their integration with real-world stimulation hardware is essential. However, this integration is complicated by the diversity of stimulation devices and experimental environments. Most models are built on idealized assumptions, such as standardized equipment configurations and accurate parameter settings, which may not reflect real-world conditions. In practice, factors such as head curvature and subject motion during experiments can result in uneven contact pressure between electrodes and the scalp. This may displace electrodes or alter contact quality, ultimately distorting the electric field distribution and reducing stimulation efficacy [iyer2005safety]. Although methods have been developed for real-time electrode localization [Clausner2017PhotogrammetryBasedHD, JaffeDax2019VideobasedMR], current tES frameworks generally lack the capability to dynamically account for such variabilities and adapt the model accordingly. Recent studies have begun to explore more adaptive and online modeling strategies to address this limitation [liu2024modelling, liang2022online]. Additionally, variability in electrode properties—such as size, material, and surface texture—can significantly influence stimulation outcomes [lopez2014dry, ng2022multi]. Despite its importance, incorporating these properties into optimization algorithms remains challenging, largely due to the absence of well-defined models to quantify their effects. While electrode size has been shown to play a critical role, no universally accepted equations currently exist to capture the influence of other electrode characteristics, such as material composition and geometry.
@@
The analysis of brain signals during tES presents a promising direction, particularly with the emergence of tES devices that support simultaneous neural signal acquisition, thereby enabling the development of closed-loop systems. Among these, EEG is frequently employed alongside tES due to the compatibility of their underlying principles. However, recording EEG during ongoing tES poses significant technical challenges. The stimulation introduces substantial artifacts, often several orders of magnitude larger than the neural signals of interest, and may even saturate EEG amplifiers [liu2018immediate, noury2017phase]. In addition to electrical artifacts, physiological noise from sources such as cardiac and respiratory activity further contaminates the EEG recordings [noury2017phase, noury2016physiological]. As a result, the signal-to-noise ratio (SNR) is drastically reduced, making direct interpretation of EEG during tES infeasible. To mitigate this issue, denoising techniques have been proposed. Several offline denoising algorithms have been developed for tES-EEG data [haslacher2021stimulation, kohli2015removal, vosskuhl2020signal], and more recently, real-time approaches have emerged to facilitate online monitoring of brain activity. For example, Guarnieri et al. introduced an adaptive spatial filtering method, known as alternating current regression (AC-REG), specifically designed to attenuate tACS-induced artifacts in real time [guarnieri2020computationally]. Despite these advances, major challenges remain. The most important of these is the lack of ground truth data, which makes it difficult to verify whether the denoised EEG accurately reflects true neural activity or still contains residual artifacts. Moreover, since tES itself modulates brain activity, the denoising process risks removing not only artifacts but also genuine stimulation-induced neural changes. As a result, validating the effectiveness and fidelity of denoised signals remains an open and actively investigated issue [kohli2020machine].
@@
The frequency of the injected current is another critical factor that warrants careful consideration in computational modeling. Contemporary models predominantly emphasize the amplitude of stimulation and the resulting electric field distribution, often overlooking the significant influence of current frequency. In the case of tACS, which typically operates at low frequencies, the quasi-static approximation is commonly employed. This approximation models the electric field at peak current amplitudes, under the assumption that temporal variations in current intensity only scale the magnitude of the electric field without affecting its spatial distribution. Nevertheless, an increasing number of studies have reported that frequency and phase can substantially affect both the distribution and efficacy of stimulation [saturnino2017target, wang2021influence]. These effects become especially pronounced at the higher carrier frequencies used in tTIS. For instance, Liu et al. observed a slight but measurable decrease in the field magnitude at temporal interference frequencies exceeding 2 kHz, which may be attributable to frequency-dependent changes in tissue conductivity [Liu2024TemporalIS, Gabriel1996TheDP]. Beyond physical field properties, frequency also plays a critical role in shaping neural responses. Variations in stimulation frequency can significantly modulate neuronal excitability and synaptic dynamics, influencing the effectiveness of neuromodulation [hutcheon2000resonance]. Therefore, it is imperative to expand computational models to incorporate frequency-dependent physiological responses. This can be achieved by integrating large-scale concurrent neural recordings and leveraging advanced signal analysis methods. Such data-driven approaches are essential for developing predictive models that elucidate the complex relationship between stimulation frequency and brain function.
@@
Additionally, the orientation of the electric field is pivotal, as it can dictate the excitatory or inhibitory responses of neurons. Especially for tTIS, which regulated brain by two pairs of electrodes, a given field orientation can significantly enhance the focality of the envelope field. While optimization algorithms have been developed to refine this aspect, there remains a gap in our understanding of how to precisely configure the field’s direction for specific targets. In the case of the cerebral cortex, a vertical orientation is commonly employed [dmochowski2011optimized]. However, for other tissues, defining a more precise target necessitates a deeper integration of neuroscience knowledge. Designing the objective function based on advanced neuroimaging could be a viable approach. For instance, fiber orientation obtained through diffusion MRI could be incorporated into the objective function.